Convert each part of ratio to percentage:
(A). 3 : 1
(B). 2 :3 : 5
(C). 1 : 4
(D). 1 : 2 : 5

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
We will be using the concepts of ratio and proportion to solve the problem also we will be using percentage concept to convert the ratio to percentage. Complete step-by-step solution - Now, we have to convert ratio to percentage. So we know that to convert a ratio to percentage, we will first convert it to fraction as x:y=xy then we multiply it by 100 and put the percentage sign. So, in (a) part we have, 3 : 1 So 3:1=31 Now we multiply it by 100. Therefore, =3×100 =300 Now, in (b) part we have a ratio as 2 : 3 : 5. So, we will let the percentage of each be 2x, 3x, 5x. Now, we know the total percentage is 100%. Therefore, 2x+3x+5x=100 10x=100 x=10 Therefore, we can write 2 : 3 : 5 as 20%, 30%, 50% by substituting the value of x=10 in 2x, 3x, 5x. Now in (c) part we have the ratio as 1 : 4. Now, we will represent this as the rational number as 14 . Now we will multiply it by 100 to obtain the percentage. Therefore, 1004=25 . So, 1 : 4 in percentage is 25%. Now in (d) part we have the ratio as 1 : 2 : 5. So, we will let the percentage of each be 1x, 2x, 5x. Now we know the total percentage is 100%. Therefore, 1x+2x+5x=100 7x=100 x=1007 Therefore, we can write 1 : 2 : 5 as 1007 by substituting the value of x=1007 in 1x, 2x, 5x respectively as 2007 and 5007. Note: To solve these types of questions one must know the concepts of ratio and proportion. One must know that if ratio is x:y then in rational can be represented as xy and to find its percentage we multiply it by 100 as xy×100.